Evaluate and tabulate the values of y for the value(s) given.
y = 2x3, for x = -3, 0, 1, 4

Answer by palanisamy(496)

Given, y = 2x^3
when, x=-3, y=2*(-3)^3=2*(-27) = -54
when, x=0, y=2*(0)^3=2*(0) = 0
when, x=1, y=2*(1)^3=2*(1) = 2
when, x=4, y=2*(4)^3=2*(64) = 128